OpenStudy (anonymous):

15. The rectangular room is given by the trinomial x^2+4x-12. What are the possible dimensions of the rectangle? Use factoring. (1 point) (x-6) and (x-2) (x+6) and (x+2) (x-6) and (x+2) (x+6) and (x-2) 17. What is the factored form of the following expression? X^2+2xy-48y^2 (1 point) a. (x+6y)(x-8y) b. (x-6y)(x-8y) c. (x+6y)(x+8y) d. (x-6y)(x+8y) 18. The area of a rectangular art studio is given by the trinomial 24x^2-4x-8. What are the possible dimensions of the studio? Use factoring. a. (3x-2) and (8x-4) b. (3x-2) and (8x+4) (3x-2) and (8x-4) (3x+2) and (8x-4) @dpasingh

OpenStudy (anonymous):

Just those @dpasingh , and thanks so much. :))

OpenStudy (anonymous):

x^2+4x-12 = x^2+6x-2x-12 = x(x+6)-2(x+6) = (x+6)(x-2) are the rquired dimensions.

OpenStudy (anonymous):

Remaining will solve short while after having my dinner...

OpenStudy (anonymous):

\[\huge x^2+2xy-48y^2 = x^2+8xy-6xy-48y^2\] \[\huge = x(x+8y)-6y(x+8y) = (x+8y)(x-6y)\] Hence (x+8y)(x-6y) are the required factors of the give expression So option d. (x-6y)(x+8y) is the correct answer... @anon99

OpenStudy (anonymous):

Thanks and the last one? @dpasingh

OpenStudy (anonymous):

\[\huge 24x^2-4x-8 = 4(6x^2-x-2)\] \[\huge = 4(6x^2-4x+3x-2)\] \[\huge = 4[2x(3x-2)+1(3x-2)] \] \[\huge = 4(3x-2)(2x+1)\] \[\huge = (3x-2)(8x+4)\] Thus (3x-2)(8x+4) are the required dimensions of the given rectangular studio.. @anon99

OpenStudy (anonymous):

THANKS SO MUCH! @dpasingh

OpenStudy (coledrf):

did you get the answers to this test? @anon99